# An Analysis of Hyper-Parameter Optimization Methods for Retrieval Augmented Generation

### v2

@@ Abstract
Finding the optimal Retrieval-Augmented Generation (RAG) configuration for a given use case can be complex and expensive. Motivated by this challenge, frameworks for RAG hyper-parameter optimization (HPO) have recently emerged, yet their effectiveness has not been rigorously benchmarked. To address this gap, we present a comprehensive study involving 5 HPO algorithms over 5 datasets from diverse domains, including a new one collected for this work on real-world product documentation. Our study explores the largest HPO search space considered to date, with three evaluation metrics as optimization targets. Analysis of the results shows that RAG HPO can be done efficiently, either greedily or with random search, and that it significantly boosts RAG performance for all datasets. For greedy HPO approaches, we show that optimizing model selection first is preferable to the prevalent practice of optimizing according to RAG pipeline order.
An Analysis of Hyper-Parameter Optimization Methods
for Retrieval Augmented Generation
Matan Orbach, Ohad Eytan, Benjamin Sznajder, Ariel Gera, Odellia Boni, Yoav Kantor, Gal Bloch, Omri Levy, Hadas Abraham, Nitzan Barzilay, Eyal Shnarch, Michael E. Factor, Shila Ofek-Koifman, Paula Ta-Shma, Assaf Toledo IBM Research matano@il.ibm.com
@@ 1 Introduction
In the Retrieval-Augmented Generation (RAG) paradigm, a generative LLM answers user questions using a retrieval system which provides relevant context from a corpus of documents (Lewis et al., 2020; Huang and Huang, 2024; Gao et al., 2024; Wang et al., 2024b). By relying on a dedicated retrieval component, RAG solutions focus LLMs on grounded data, reducing the likelihood of dependence on irrelevant preexisting knowledge.
@@
Furthermore, evaluating even a single RAG configuration is costly in terms of time and funds: the embedding step as part of corpus indexing is compute-intensive; generating answers using LLMs is also a demanding task, especially for large benchmarks; and evaluation with LLMaaJ adds another costly round of inference. Therefore, exploring and evaluating the full exponential search space of RAG parameters is prohibitively expensive; and at the same time, suboptimal parameter choices may significantly harm result quality.
One approach for addressing these issues is to use automated hyper-parameter optimization (HPO) for RAG (Fu et al., 2024; Kim et al., 2024): conducting experiments on a limited set of RAG configurations, and recommending a configuration based on the results. Experiments can be performed in various ways, ranging from established HPO algorithms to a naive random sampling of configurations to test. Importantly, although HPO frameworks are gaining ground in RAG use-cases, their effectiveness has not been rigorously tested.
@@
Figure 1: We study hyper parameter optimization over a RAG pipeline with $5$ parameters. The explored search space includes $162$ RAG configurations formed from combinations of the depicted hyper parameters.
@@
In this work we study the question of how well HPO algorithms work in practice for standard RAG pipelines comprising retrieval and generation (see Figure 1). Our experiments explore the largest RAG search space considered to date, consisting of $162$ possible RAG configurations, derived from five parameters: chunk size and overlap, which control how documents are split, the embedding model used to encode chunks in a vector database, the number of retrieved chunks included as context when answering a question, and the generative model that produces the answer. This search space is explored with $5$ HPO algorithms: Tree-Structured Parzen Estimators (TPE) (Watanabe, 2023), $3$ greedy algorithms, and random search. Our evaluation also includes grid search over the entire search space, enabling comparison to the best possible result. The study of a larger search space induced by more complex RAG pipelines with additional parameters (e.g. agentic pipelines, adding rerankers, etc.) is computationally demanding for grid search and deferred to future work.
Our experimental setup follows the real-world scenario where a dataset is provided up front for experimentation and unseen user questions arrive after deployment. We therefore select RAG configurations based on their performance on a development set, and report their results on held out test data. To represent real use cases, we employ a diverse set of $4$ existing RAG question-answering datasets from different domains: arXiv articles on ML (Eibich et al., 2024), Biomedical (Krithara et al., 2023), Wikipedia (Rosenthal et al., 2024; Smith et al., 2008). In addition, we experiment with a new fifth dataset over the domain of product documentation, which was collected and open-sourced as part of this work. Lastly, as different applications optimize according to different dimensions (e.g. answer correctness or faithfulness), our study also takes into account different RAG objectives implemented in various ways.
@@
Our main contributions are: (i) benchmark results over the largest RAG search space to-date, showing that RAG HPO can be done efficiently, either greedily or with simple random search, and that it significantly boosts RAG performance for all datasets; (ii) a detailed analysis of the results, exploring the connections between the optimized parameters, the dataset and the optimization objective. For greedy HPO approaches, we show that the order in which the parameters are optimized is of great importance; (iii) a new open-source RAG dataset over enterprise product documentation. [^1] https://huggingface.co/datasets/ibm-research/watsonxDocsQA
@@ 2 Related Work
Prior work on HPO for RAG includes AutoRAG-HP Fu et al. (2024), which explores $75$ RAG configurations induced from $3$ parameters. In comparison, our search space spans $162$ configurations, and adds the choice of a generative model, which we show to be the most dominant RAG parameter (see §6). Importantly, Fu et al. (2024) focus on a setting with online feedback from users, and thus explore an online algorithm that iteratively updates the reward for the various RAG parameters based on small batches of queries. In contrast, our study involves offline HPO procedures which take full benchmarks as input. Also, Fu et al. (2024) report results on the same set they optimize on, hiding potential over-fitting, while our reports are on a held-out test set, simulating real-world performance on unseen user queries.
Other work focuses on manual tuning of RAG hyper-parameters, for example CRUD-RAG Lyu et al. (2024) and Wang et al. (2024c). While a manual recipe for finding a good RAG configuration is advantageous, automation is essential to generalize to new use cases and domains, and to handle continually evolving datasets and models.
More recently, RAGEval Zhu et al. (2024) uses synthetic benchmark generation techniques to generate the DragonBall “omni” benchmark, which encompasses data from finance, law and medical texts in English and Chinese. They found that certain hyper-parameters are strongly affected by dataset, such as the optimal chunk size and the number of retrieved chunks. This provides further motivation for automated RAG HPO.
Within the open-source community, multiple tools offer out-of-the-box HPO algorithms for RAG. One is the AutoRAG framework (Kim et al., 2024), which takes a greedy approach, optimizing one RAG parameter at a time based on a sequential pipeline order. [^2] The AutoRAG repository. Another is RAGBuilder which uses TPE. [^3] The RAGBuilder repository. RAG-centric libraries such as LlamaIndex [^4] The LlamaIndex repository. also support HPO by integrating with similar open source frameworks like RayTune Liaw et al. (2018), optuna Akiba et al. (2019) and hyperopt Bergstra et al. (2013).
Missing from all the above works is a systematic study that shows how well these algorithms work in practice, and compares them with each other over multiple datasets. That is the main focus of our work. The greedy algorithms we explore in §4 are similar in spirit to the greedy algorithm used by Kim et al. (2024), and the TPE algorithm we evaluate is the same one employed by RAGBuilder.
@@ 3 Experimental Setup
In our explored RAG pipeline (see Figure 1), processing starts with chunking the input documents into smaller chunks, based on two predefined parameters: the size of each chunk in tokens, and the overlap between consecutive chunks. This is followed by representing each chunk with a dense vector created by an embedding model – our third parameter. The vectors are stored in a vector-database, [^5] Milvus (Wang et al., 2021) with default settings; index type is HNSW Malkov and Yashunin (2018). alongside their original text. Upon receiving a query, it is embedded too, and the Top-K ($K$ being our fourth parameter) relevant chunks are retrieved (retrieval). Lastly, a prompt containing the query and the retrieved chunks is passed to a generative model (our fifth parameter) to create an answer (generation). [^6] Generation details are in Appendix F.
3.1 Search space
The specific values considered for each of our $5$ hyper-parameters are described in Table 1. In total, the search space has $3*2*3*3*3=162$ possible RAG configurations. Among them, $18$ ($3*2*2$) relate to data indexing (chunking and embedding). The values chosen for the chunk size, overlap, and top-k reflect common practices. Popular open source models were selected as options for the embedding and generative models.
@@ 3.1 Search space
Chunk Overlap (Tokens) Chunking 0%, 25%
Embedding Model Embedding multilingual-e5-large Wang et al. (2024a) bge-large-en-v1.5 Xiao et al. (2023) granite-embedding-125M-english IBM (2024)
@@
Generative Model Generation Llama-3.1-8B-Instruct AI (2024) Mistral-Nemo-Instruct-2407 AI and NVIDIA (2024) Granite-3.1-8B-instruct Granite Team (2024)
@@
Our experiments involve performing a full grid search – i.e., evaluating all possible configurations – in order to infer an upper bound for the optimization strategies. Hence, due to computational constraints we avoid using an even larger search space.
3.2 Datasets
AIArxiv
This dataset was derived from the ARAGOG benchmark (Eibich et al., 2024) of technical QA pairs over a corpus of machine learning papers from ArXiv. [^7] AIArxiv is available here. As gold documents are not annotated in ARAGOG dataset, we added such labels where they could be found, obtaining $71$ answerable QA pairs out of $107$ in the original benchmark.
BioASQ
(Krithara et al., 2023) A subset of the BioASQ Challenge train set. [^8] BioASQ is available here. Its corpus contains $40200$ passages extracted from clinical case reports. The corresponding benchmark of $4200$ QA pairs includes multiple gold documents per question.
MiniWiki
A benchmark of $918$ QA pairs over Wikipedia derived from Smith et al. (2008). [^9] MiniWiki is available here. The contents are mostly factoid questions with short answers. This dataset has no gold document labels.
ClapNQ
Rosenthal et al. (2024) A subset of the Natural Questions (NQ) dataset (Kwiatkowski et al., 2019) on Wikipedia pages, of questions that have long answers. The original benchmark contains both answerable and unanswerable questions. For our analysis we consider only the former. ClapNQ dataset consists of $178890$ sub-documents generated from $4293$ pages. These sub-documents constitute the input to the pipeline.
WatsonxQA
(ProductDocs) A new open source dataset and benchmark based on enterprise product documentation. It consists of $5534$ sub-documents generated from $1144$ pages of HTML product documentation. As with the ClapNQ dataset, the sub-documents serve as input for the pipeline. The benchmark includes $75$ QA pairs and gold document labels, of which $25$ were generated by two subject matter experts, and the rest were generated synthetically and then manually filtered. All QA pairs were additionally reviewed by two of the authors, ensuring high data quality. Further details and examples from the datasets are in Appendix A.1.
Overall, these datasets exhibit variability in many aspects. They represent diverse domains – research papers, biomedical documents, wikipedia pages and enterprise data. They also vary in question and answer lengths; for example, MiniWiki has relatively short answers, while ClapNQ was purposely built with long gold answers. Corpus sizes also vary, representing real-world retrieval scenarios over small or large sets of documents.
Every benchmark was split into development and test sets. To keep computations tractable, the number of questions in the large benchmarks (BioASQ and ClapNQ) was limited to $1000$ for development and $150$ for test. Table 2 lists the corpora and benchmark sizes and domains.
@@ 3.2 Datasets
Table 2: Properties of the RAG Datasets in our experiments, including the number of documents within the corpus and QA pairs within the #Dev and #Test sets.
3.3 Metrics
The following metrics were used in our experiments: Retrieval quality was measured using context correctness with Mean Reciprocal Rank Voorhees and Tice (2000).101010 Note that as labeling is at the document level, any chunk from the gold document is considered as a correct prediction, even if it does not include the answer to the question. Answer faithfulness (Lexical-FF) measures whether a generated answer remained faithful to the retrieved contexts, with lexical token precision. Answer correctness compares generated and gold answers, assessing overall pipeline quality, and is measured in two ways: First, a fast, lexical, implementation based on token recall (Lexical-AC), which provides a good balance between speed and quality Adlakha et al. (2024). Second, a standard LLM-as-a-Judge answer correctness (LLMaaJ-AC) implementation from the RAGAS library Es et al. (2023), with GPT4o-mini OpenAI (2024) as its backbone. This implementation performs $3$ calls to the LLM on every invocation, making it much slower and more expensive than the lexical variant.
4 HPO Algorithms
An HPO algorithm takes as input a RAG search space, a dataset (corpus + benchmark), and one or more evaluation metrics, of which one is defined as the optimization objective. Its goal is to find the best-performing RAG configuration on the dataset in terms of the objective metric.
The HPO algorithms in our experiments are iterative. At each iteration, the algorithm looks at all previously explored RAG configurations and their scores, and chooses one unexplored configuration to try next. Simulating a limited exploration budget, after a fixed number of iterations the algorithm stops and the best performing configuration is returned. A good HPO algorithm identifies a top-performing configuration with minimal iterations.
We consider two types of HPO algorithms. The first are standard HPO algorithms not specifically tailored to RAG. The other type required some knowledge of the optimized RAG pipeline and its components. All optimize answer correctness (see §3.3) unless explicitly stated otherwise.
Our first standard HPO algorithm is the Tree-Structured Parzen Estimators (TPE) algorithm Watanabe (2023), using an implementation from hyperopt Bergstra et al. (2013). It was configured with $5$ random bootstrapping iterations and otherwise the default settings. The second standard algorithm (Random) samples an unexplored random RAG configuration at each iteration, while ignoring results from previous iterations.
The other type of algorithms take as input a list of search space parameters ordered by their presumed impact on the RAG pipeline. These algorithms take a greedy approach: they iterate the list, optimizing one parameter at a time, assuming that optimizing more impactful ones first leads to quick identification of a good configuration. When optimizing a parameter $p$ the algorithm assumes that values for preceding parameters were already determined. The parameter $p$ itself is evaluated with all possible values, while the values of all following parameters are chosen randomly. The value yielding the best objective score is picked for $p$, and the algorithm continues to the next hyper-parameter.
The greedy algorithms differ solely in the order in which the hyper-parameters are optimized. One option (Greedy-M) is to order the generative and embedding models first, assuming they are more important: Generative Model, Embedding Model, Chunk Size, Chunk Overlap, Top-k. A second prevalent one (Greedy-R) is to optimize by the order of the RAG pipeline itself, first the retrieval part (still with the model first), followed similarly by the generation part: Embedding Model, Chunk Size, Chunk Overlap, Generative Model and Top-k. A third option (also prevalent) (Greedy-R-CC) uses the same order, yet optimizes the first $3$ retrieval-related parameters according to the score of a context correctness metric evaluated on the produced index, thus avoiding any LLM-based inference until all retrieval parameters are chosen. After that, the algorithm optimizes the remaining parameters with answer correctness.
5 Results
(a) LLMaaJ-AC
(b) Lexical-AC
Figure 2: Per-iteration performance of all HPO algorithms on the test sets of five datasets, optimizing answer correctness computed with an LLMaaJ metric (a) and a lexical metric (b). The dashed black lines denote the best achievable performance on each test set.
Experiments Description
The experiments consider the search space of §3.1 and all described datasets in §3.2. The optimization objective was answer correctness, implemented with either LLMaaJ-AC or Lexical-AC, or answer faithfulness with Lexical-FF (§3.3).
Our experimental setting was chosen to simulate a realistic use case: an HPO algorithm is executed over a benchmark (the development set) and the best RAG configuration is chosen for deployment; the deployed configuration is expected to generalize to a set of unseen questions – simulated here via a test set.
Specifically, every HPO algorithm ran on each development set for $10$ iterations. After each iteration, the best configuration on the development set was applied to the test set. This allows a per-iteration progress analysis (see below). Since all algorithms involve a random element, each run was repeated with $10$ different seeds.
To establish upper bounds on the expected performance, we also ran a full grid search over the development and test sets. These upper bounds allow us to assess to what extent the different algorithms approach the optimal result. This entails evaluating all possible $162$ configurations, including $18$ different indexes, on all $5$ datasets.
Main Results
Figure 2 depicts the per-iteration performance of the HPO methods on the test sets, when optimizing for answer correctness. Results for answer faithfulness are of the same trend, and are therefore deferred to Appendix §B. In each plot, the best achievable performance, obtained by running a full grid search on the test set and selecting the best configuration, is shown with a black dashed line.
For all algorithms, the plots demonstrate that it is enough to explore around $10$ configurations to obtain performance that is comparable to exploring all $162$ available configurations with a naive grid search (solid red lines in Figure 2). This strong result holds for both evaluation metrics.
In some cases there is a substantial gap between performance of the best achievable configuration on the test test (dashed black lines), and the performance obtained when taking the best configuration on the development set, and applying it to the test (red lines). This is an inherent limitation of any generalization scenario. All algorithms, being exposed only to the development set, can only aspire to performance close to the red lines on the test, and indeed the results all show they converge to roughly that performance.
Focusing on the greedy algorithms, the results suggest that the order in which the parameters are optimized is of great importance. Specifically, methods that start by optimizing retrieval-related parameters (Greedy-R and Greedy-R-CC) require more iterations to reach good configurations.
The naive option of random sampling also finds a good RAG configuration after a small number of iterations. The complex TPE algorithm was found to be roughly equivalent in quality to random choice in our setup. This presumably stems from the size of our search space. In even larger search spaces (for example, with continuous parameters) further exploration of TPE may prove valuable.
Figure 3: Per-iteration LLMaaJ-AC performance on the test sets of two largest datasets, for HPO algorithms optimized using the full development data (solid lines) or a sample of it (dotted). The dashed black lines show the best achievable test performance.
Development Set Sampling
The above results were produced with each RAG pattern evaluated on the whole development set. While simple, this option is costly for large datasets. Research in other domains has suggested that limiting evaluation benchmarks by random sampling improves efficiency without compromising evaluation quality Perlitz et al. (2024); Polo et al. (2024). Thus, we experiment with HPO that uses a sample of the development set. To our knowledge, this is the first study of sampling in the context of RAG HPO.
To match the RAG setting, we sample both the benchmark and the underlying corpus, to reduce both inference and embedding costs. Specifically, focusing on the larger datasets of BioASQ and ClapNQ, 10% of the development set QA pairs were sampled along with their corresponding gold documents – those containing the answers to the sampled questions. To simulate realistic retrieval scenarios we add “noise” – documents not containing an answer to any of the sampled questions – at a ratio of $9$ such documents per gold document. This process results in $100$ sampled benchmark questions per dataset. The sampled corpora comprise $1$K (i.e. $1000$) documents for ClapNQ (out of $178$K), and $10$K for BioASQ (out of $40$K). [^11] BioASQ has multiple gold documents per question, which yields more sampled documents.
Following sampling, we repeat the experiments using the best HPO methods: Random, TPE, and Greedy-M. Figure 3 depicts the results when optimizing for LLMaaJ-AC, comparing all three methods when applied to the full (solid lines) or sampled (dotted) development sets (similar Lexical-AC results are in Appendix B). For BioASQ the results remain about the same, while for ClapNQ sampling leads to identifying a suboptimal configuration after all $10$ iterations, most notably for TPE and Random.
In terms of cost, for both datasets, performing inference for a given configuration is $10$x cheaper (as $10$% of the questions are used). Indexing is also far cheaper, since the corpus is decreased by $4$x for BioASQ and $178$x for ClapNQ.
Overall, while sampling the development set is computationally promising, our mixed results suggest it remains an open challenge – i.e., an appropriate sampling approach must be used or else quality is degraded. We leave this for further work.
6 Analysis
We analyze three aspects of our experiments: the impact of the chosen evaluation metric, the impact of retrieval quality on answer quality, and cost.
(a) LLMaaJ-AC
(b) Lexical-AC
Figure 4: The effect of chosen optimization metric on the generative model within the best RAG configuration.
Impact of Metric Choice
According to the results of Greedy-M in Figure 2, performance is boosted significantly when optimizing the generative model parameter first, suggesting its importance. We therefore further examine the best performing model in each experiment.
Figure 4 depicts per-dataset maximal answer correctness scores for each generative model (maximum of all $162/3=54$ configurations the model participates in). Interestingly, the best RAG configuration for all datasets involves Llama when optimizing by LLMaaJ-AC, and Granite or Mistral when optimizing by Lexical-AC. This difference stems from the metrics, as Lexical-AC is recall-oriented while LLMaaJ-AC accounts for both precision and recall. This highlights the importance of selecting a target metric that correctly reflects the use case preferences, as optimal configurations may vary for different target metrics.
Figure 5: The effect of retrieval quality (context correctness, x-axis) on generation quality (answer correctness, y-axis). Each point denotes one of the $162$ RAG configurations in the search space.
Impact of Retrieval Quality
Figure 5 depicts the relationship between context correctness and answer correctness. The two measures are correlated – when retrieval quality is low, so is answer quality. Also evident is the influence of the generative model choice on answer quality.
The impact of context correctness on answer correctness diminishes as context correctness increases, as implied by the line slopes in Figure 5 being steeper at low context correctness values. This behavior can be attributed to the fact that several retrieved chunks are utilized during generation. When context correctness is low, there is a high probability that none of the chunks are relevant, and the generated answer will therefore be incorrect. Conversely, when at least one relevant chunk is retrieved, the quality of the answer will largely rely on generation hyper-parameters.
Cost Considerations
The cost of each HPO algorithm was tracked by counting the number of tokens embedded when creating each index, and the number of input and output tokens used in generation. These counts are accumulated per-algorithm over all explored configurations till some specific iteration. Indices that are reused in multiple explored configurations are accounted for once. Detailed plots of these counts are in Appendix C.
Overall, the generation cost was about the same for all algorithms and datasets. The embedding cost, being a function of the number of different indices created by each algorithm, behaved similarly across datasets with some variations between algorithms. In this regard, both TPE and Random have no built-in preference to reuse already constructed indices. In contrast, both Greedy-R and Greedy-R-CC create a new index in each iteration till they have optimized all index-related parameters, and are therefore initially costly. Greedy-M is initially cheap, since starts by optimizing the generative model parameter using just one index. It is therefore preferable if cost is crucial and the limit on the number of explored configuration is low.
7 Conclusion
Our work presented a comprehensive study of multiple RAG HPO algorithms, exploring their performance over datasets from various domains and using multiple metrics, and demonstrated that RAG HPO can boost performance significantly. Compared to an arbitrary choice of RAG configurations, $10$ HPO iterations improve performance by as much as $20$% (see Figure 2, comparing performance at iterations $1$ and $10$). Since this holds for simple RAG pipelines, the effect on more complex (or agentic) RAG pipelines with many more hyper-parameters is likely to be even greater, a subject for further work.
We show that RAG HPO can be performed efficiently, exploring only a fraction of all possible RAG configurations induced by the hyper-parameters. Even without prior knowledge of the RAG pipeline architecture, iterative random sampling bounded by a fixed number of iterations can suffice. We also show that a greedy HPO approach that optimizes models first is preferable to the prevalent practice of optimizing sequentially according to pipeline order.
Our results highlight the importance of the chosen optimization objective, as different choices of metrics correspond to different optimal RAG configurations. Jointly optimizing multiple metrics is an interesting direction for future work.
We take initial steps in improving the efficiency of RAG HPO, by sampling the development set, and plan to further explore this in future work. We also contribute a new RAG dataset, of enterprise product documentation, for use by the community.
Limitations
We limited the size of our search space to enable a full grid search on all datasets in a reasonable time and cost. Further work could explore a larger search space without comparing to a full grid search.
Our experiments optimized a simple linear RAG workflow comprising retrieval followed by generation. Additional RAG approaches may introduce additional hyper-parameters to the mix, such as filters, rerankers, answer reviewers, HyDE etc. Moreover, our prompts were fixed per generation model and we did not explore prompt optimization.
Our analysis considers only English textual datasets with answerable questions. The comparison of performance for multi-modal, multi-lingual datasets that include also unanswerable questions is left for future work.
It should be noted that our method for retrieval quality evaluation is not accurate: any chunk from the gold document is considered a correct prediction even if it does not include the answer to the question.
Appendix A Datasets
Example questions per dataset are listed in Table 3.
A.1 WatsonxQA additional details
As stated in the body of the paper, the WatsonxQA benchmark includes $75$ QA pairs and gold document labels, of which $50$ were generated synthetically. The benchmark contains five fields: a question, a gold answer and, for computing context relevance metrics, the gold passage id and its content. Some example benchmark entries are given in Figure 12. Synthetic generation was operated using falcon-180b model, and then manually filtered and reviewed for quality - the methodology we used is detailed in Yehudai et al. (2024).
The prompt used for the synthetic generation is detailed in  Figure 13.
Dataset Example Query

### v3

@@ Abstract
Optimizing Retrieval-Augmented Generation (RAG) configurations for specific tasks is a complex and resource-intensive challenge. Motivated by this challenge, frameworks for RAG hyper-parameter optimization (HPO) have recently emerged, yet their effectiveness has not been rigorously benchmarked. To fill this gap, we present a comprehensive study involving five HPO algorithms over five datasets from diverse domains, including a newly curated real-world product documentation dataset. Our study explores the largest RAG HPO search space to date that includes full grid-search evaluations, and uses three evaluation metrics as optimization targets. Analysis of the results shows that RAG HPO can be done efficiently, either greedily or with random search, and that it significantly boosts RAG performance for all datasets. For greedy HPO approaches, we show that optimizing model selection first is preferable to the common practice of following the RAG pipeline order during optimization.
@@ 1 Introduction
In the Retrieval-Augmented Generation (RAG) paradigm, a generative LLM answers user questions using a retrieval system which provides relevant context from a corpus of documents (rag1; rag3; rag2; rag4). By relying on a dedicated retrieval component, RAG solutions focus LLMs on grounded data, reducing the likelihood of dependence on irrelevant preexisting knowledge.
@@
Furthermore, evaluating even a single RAG configuration is costly in terms of time and funds: the embedding step as part of corpus indexing is compute-intensive; generating answers using LLMs is also a demanding task, especially for large benchmarks; and evaluation with LLM-as-a-Judge (LLMaaJ) adds another costly round of inference. As a result, exhaustively exploring the exponential search space of RAG parameters is prohibitively expensive. At the same time, suboptimal parameter choices may significantly harm result quality.
A promising approach to this challenge is hyper-parameter optimization (HPO) for RAG (fu-etal-2024-autorag; kim2024autorag), which aims to identify high-performing configurations by systematically evaluating a subset of the search space. Existing methods range from established HPO algorithms to simple random sampling. Importantly, despite growing interest, the effectiveness of HPO in realistic RAG scenarios has not been rigorously tested.
@@
Figure 1: We study hyper parameter optimization over a RAG pipeline with 55 parameters. The explored search space includes $162$ RAG configurations formed from combinations of the depicted hyper parameters.
Our work addresses this gap through a comprehensive study of HPO for RAG in a setup that mirrors real-world usage, where a dataset is provided up front for experimentation, and unseen queries arrive after deployment. The best RAG configuration is therefore selected by development set performance and evaluated on held out test data. To the best of our knowledge, this is the first study that considers RAG HPO in such a setup.
The scope of our experiments includes multiple datasets, evaluation metrics and algorithms. To represent diverse use cases, we evaluate across several domains: scientific (eibich2024aragog), biomedical (krithara2023bioasq), Wikipedia (rosenthal2024clapnq; smith2008question) and a newly curated enterprise product documentation dataset that we open-source as part of this work. [^1] http://huggingface.co/datasets/ibm-research/watsonxDocsQA Additionally, recognizing that different applications prioritize distinct performance metrics – such as answer correctness or faithfulness – our study evaluates multiple RAG optimization objectives, implemented through two alternative approaches: lexical overlap metrics and LLMaaJ.
Our evaluation compares five HPO algorithms: Tree-Structured Parzen Estimators (TPE) (tpe), three greedy variants, and random search. Also included are baseline grid search results for both the development and test sets, over the full search space. [^2] The grid results are at https://github.com/IBM/rag-hpo-bench. These are crucial for establishing upper bounds on HPO performance for RAG, and understanding its generalization capability.
While the inclusion of grid search enables a comparison to the best possible result, it poses computational constraints on the size of the explored search space. Nonetheless, our search space is the largest considered for RAG HPO to date while still including a comparison to full grid search.
@@
The search space comprises of $162$ RAG configurations derived from five core RAG parameters (see Figure 1): chunk size and overlap, which control how documents are split, the embedding model used to encode chunks in a vector database, the number of retrieved chunks included as context when answering a question, and the generative model that produces the answer. Exploring an expanded search space introduced by more complex RAG pipelines, such as agentic workflows, poses significant computational challenges for exhaustive grid search. Consequently, this aspect is deferred to future work.
Our evaluation addresses multiple aspects of HPO for RAG, including the convergence properties of the algorithms, the impact of the objective metric on the best configuration, and an analysis of HPO overall cost. We also explore reducing that cost through development set sampling.
@@
In summary, our main contributions are as follows: (i) comprehensive benchmarking of RAG HPO in a realistic generalization setup, over the largest search space with full grid-search evaluations to-date, showing that RAG HPO can be done efficiently, either greedily or with simple random search, and that it significantly boosts RAG performance for all datasets; (ii) a detailed analysis of the results, exploring the connections between the optimized parameters, the dataset and the optimization objective. For greedy HPO approaches, we show that the order in which the parameters are optimized is of great importance; (iii) new open-source resources: the full grid search results of our experiments, and an enterprise product documentation RAG dataset.
@@ 2 Related Work
Within the open-source community, several tools offer out-of-the-box HPO algorithms for RAG. AutoRAG (kim2024autorag) adopts a greedy approach, optimizing one RAG parameter at a time following the sequential pipeline order. [^3] https://github.com/Marker-Inc-Korea/AutoRAG RAGBuilder employs TPE for HPO. [^4] https://github.com/KruxAI/ragbuilder Additionaly, RAG-centric libraries such as LlamaIndex [^5] https://www.llamaindex.ai/ support HPO by integrating general-purpose optimization frameworks like RayTune (liaw2018tune), optuna (optuna) and hyperopt (hyperopt).
Other works investigate the impact of RAG hyper-parameters without automated optimization. For example, crud-rag and wang2024searching focus on manual tuning of RAG hyper-parameters, while ragbench evaluates multiple configurations via grid search. Their studies highlight the critical role of hyper-parameter tuning in RAG and motivate the need for automated RAG HPO.
Another interesting line of work (fu-etal-2024-autorag) is motivated by a setting of online feedback from users. It describes an online HPO algorithm that iteratively updates the reward for the various RAG parameters based on small batches of queries. In contrast, our work addresses offline HPO, where optimization is performed on full benchmark datasets prior to best configuration deployment.
More recently, barker2025multi_objective_optimization introduced multi-objective HPO for RAG. The studied methods select a set of RAG configuration deemed Pareto-optimal by multiple metrics. Choosing the best configuration from this set remains an open challenge. Our study instead focuses on single-objective HPO returning a single configuration as output.
Despite these efforts, still missing is a systematic evaluation of HPO algorithms across diverse datasets under realistic conditions where optimized configurations are tested on held-out sets. This gap is the primary focus of our work. Building upon existing approaches, our evaluation prioritizes HPO algorithms already considered in the context of RAG, over introducing alternatives such as BOHB (falkner2018bohb) or SMAC (lindauer2022smac3). Specifically, the greedy algorithms we explore resemble those used by kim2024autorag, and the TPE algorithm we assess is the same one used by RAGBuilder.
@@ 3 Experimental Setup
Search space
@@
Chunk Overlap (% Tokens) Chunking 0%, 25%
Embedding Model Embedding multilingual-e5-large (multilingual-e5-large) bge-large-en-v1.5 (bge_embedding) granite-embedding-125M-english (ibm2024graniteembedding125m)
@@
Generative Model Generation Llama-3.1-8B-Instruct (meta2024llama3.1) Mistral-Nemo-Instruct-2407 (mistral2024nemo) Granite-3.1-8B-instruct (granite2024)
@@
In our explored RAG pipeline (see Figure 1), processing starts with chunking the input documents into smaller chunks, based on two parameters: the size of each chunk in tokens, and the overlap between consecutive chunks. This is followed by representing each chunk with a dense vector created by an embedding model – our third parameter. The vectors are stored in a vector-database, [^6] Milvus (milvus) with default settings; index type is HNSW (hnsw). alongside their original text. Upon receiving a query, the top $k$ relevant chunks are retrieved (retrieval); $k$ being our fourth parameter. Lastly, a prompt containing the query and the retrieved chunks is passed to a generative model (our fifth parameter) to create an answer (generation). Greedy decoding was used throughout all experiments. The prompts were fixed to RAG prompts tailored to each model. [^7] See Appendix G.
The specific values considered for each of the five hyper-parameters are described in Table 1. In total, the search space has 32333=1623*2*3*3*3=162 possible RAG configurations. Dividing into stages, we get $18$ (3233*2*3) different configurations for data indexing (chunking and embedding) and 99 (333*3) different configurations for answering (retrieval and generation). The values chosen for the chunk size, overlap, and top-k reflect common practices (see Appendix B for details). Popular open source models were selected as options for the embedding and generative models. We opted to focus on LLMs that are similar in size, since otherwise an obvious strategy is to simply discard the smaller less capable models from the search.
Our experiments involve performing a full grid search – i.e., evaluating all possible configurations – in order to establish upper bound baselines for the optimization strategies. Hence, due to computational constraints we avoid using an even larger search space, and choose a set of moderately sized LLMs as our generators. Still, to the best of our knowledge, our search space is the largest considered to date that compares to full grid-search.
Datasets
@@
Table 2: Properties of the RAG Datasets in our experiments: the number of documents within the corpus (#Doc), and counts of QA pairs in the #Dev and #Test sets.
AIArxiv
This dataset was derived from the ARAGOG benchmark (eibich2024aragog) of technical QA pairs over a corpus of machine learning papers from ArXiv. [^8] https://huggingface.co/datasets/jamescalam/ai-arxiv2 As gold documents are not annotated in ARAGOG dataset, we added such labels where they could be found, obtaining $71$ answerable QA pairs out of $107$ in the original benchmark.
BioASQ
(krithara2023bioasq) A subset of the BioASQ Challenge train set. [^9] https://huggingface.co/datasets/rag-datasets/rag-mini-bioasq Its corpus contains $40200$ passages extracted from clinical case reports. The corresponding benchmark of $4200$ QA pairs includes multiple gold documents per question.
MiniWiki
A benchmark of $918$ QA pairs over Wikipedia derived from smith2008question. [^10] https://huggingface.co/datasets/rag-datasets/rag-mini-wikipedia The contents are mostly factoid questions with short answers. This dataset has no gold document labels.
ClapNQ
(rosenthal2024clapnq) A subset of the Natural Questions (NQ) dataset (kwiatkowski2019natural) on Wikipedia pages, of questions that have long answers. The original benchmark contains both answerable and unanswerable questions. For our analysis we consider only the former. ClapNQ dataset consists of $178890$ passage texts generated from $4293$ pages. These passages constitute the input to the pipeline.
WatsonxQA
(ProductDocs) A new open-source dataset and benchmark based on enterprise product documentation, consisting of $5534$ passage texts created from $1144$ HTML product documentation pages. [^11] http://huggingface.co/datasets/ibm-research/watsonxDocsQA These passages serve as the RAG pipeline input. The benchmark includes $75$ QA pairs and gold document labels, of which $25$ were generated by two subject matter experts, and the rest were synthetically produced and then manually filtered. All QA pairs were additionally reviewed by two of the authors, ensuring high data quality. Further details are in Appendix A.
Overall, these datasets exhibit variability in many aspects. They represent diverse domains – research papers, biomedical documents, wikipedia pages and enterprise data (see question examples in Table 3). They also vary in question and answer lengths; for example, MiniWiki has relatively short answers, while ClapNQ was purposely built with long gold answers. Corpus sizes also vary, representing real-world retrieval scenarios over small or large sets of documents.
Every benchmark was split into development and test sets. To keep computations tractable, the number of questions in the large benchmarks (BioASQ and ClapNQ) was limited to $1000$ for development and $150$ for test. Table 2 lists the corpora benchmark sizes and domains.
Dataset Example Question